OpenStudy (anonymous):

Which of the following algebraic expressions represents the verbal expression given below? The opposite of a number, raised to the fourth power, is divided by five.

OpenStudy (anonymous):

of a number =x opposite is just - so we start with a -x

OpenStudy (anonymous):

-x is raised to the 4th power....can you tell me what it's going to look like now?

OpenStudy (anonymous):

hint: it has to do with exponents

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes that looks about right. good job

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

Wait. There is a mistake.

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

You need to be careful with the square of -x. There is a difference between \(-x^2\) and \( (-x)^2\). \( -x^2\) means take the square of x, then find its opposite. \((-x)^2 \) means find the opposite of x, then square it.

OpenStudy (anonymous):

Well it doesnt matter in this case because the exponent value is positive.

OpenStudy (mathstudent55):

Here is an example: Let's say x = 4. \(-x^2 \) means: square x: \(4^2 = 16\); now find its opposite: the opposite of 16 is -16. The answer is -16 \( (-x)^2 \) means find the opposite of x. The opposite of 4 is -4. Now square it. The square of -4 is (-4)(-4) = 16. The answer is 16. As you can see, one expression gives you -16 , and the other one gives you 16.

OpenStudy (anonymous):

I see, can you write a verbal expression for the other expression then to compare.

OpenStudy (anonymous):

so what would the anser look like?

OpenStudy (mathstudent55):

Let's take your problem from the beginning. The opposite of a number, raised to the fourth power, is divided by five. A number: x The opposite of a number: -x raised to the fourth power: \( (-x)^4 \) is divided by five. \( \dfrac{(-x)^4}{5} \) This is your final answer.

OpenStudy (mathstudent55):

Great question. \( - \dfrac{x^4}{5} \) A number raised to the 4th power is divided by five. Then the opposite is taken.

OpenStudy (anonymous):

@mathstudent55 Great job! I did it so fast I forgot about this!

OpenStudy (anonymous):

thanks both of you for the help! :)

OpenStudy (mathstudent55):

wlcm